OpenStudy (anonymous):

Consider the function f(x)=3x^2/x-2. (a)Set up the f1(x)=lim f(w)-f(x)/w-x. (b) set up the f1(x)=lim f(x+h)-f(x)/h

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

For what?

OpenStudy (helder_edwin):

(a) \[ \large f'(x)=\lim_{w\to x}\frac{f(w)-f(x)}{w-x}= \lim_{w\to x}\frac{\frac{3w^2}{w-2}-\frac{3x^2}{x-2}}{w-x} \] can u go on from this?

OpenStudy (anonymous):

Oh yeah I got that too, I have a much harder time with the other form.

OpenStudy (helder_edwin):

(b) \[ \large f'(x)=\lim_{h\to0}\frac{f(x+h)-f(x)}{h}= \lim_{h\to0}\frac{\frac{3(x+h)^2}{(x+h)-2}-\frac{3x^2}{x-2}}{h} \]

OpenStudy (helder_edwin):

can u finish computing both limits? it is quite easy.

OpenStudy (anonymous):

Yeah I think I can thanks for your help, you are a lifesaver :)

OpenStudy (helder_edwin):

u r welcome.